OpenStudy (anonymous):

How do I solve 10x=6(2x-3)

OpenStudy (anonymous):

Distribute the 6 to the (2x-3) so now its 12x-18 = 10x then subtract the 10 x from the 12 x So now its 2x-18=0 Then 2x=18 x=18/2 x=9

OpenStudy (anonymous):

distribute the six to get 12x-18 rewrite the equation 12x-18=10x subtract the 12x from the 10x rewrite:-2x=-18 then divide -18 by -2 x=9

OpenStudy (anonymous):